OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9?

OpenStudy (anonymous):

y+3=1(x-2) y=x-1 then you change to the standard form which I believe to be Ax+By=C

OpenStudy (anonymous):

wait, where did you get the first part from?

OpenStudy (anonymous):

I used point slope form it is y-y1=m(x-x1) the slope is m and since it is parallel it is going to be 1 y1=-3 and x1=2 these I got form the point you gave (2,-3)

OpenStudy (anonymous):

the answer is x-y=1?

OpenStudy (anonymous):

It appears to be so

OpenStudy (anonymous):

thanks for your explanation :)

OpenStudy (anonymous):

You're welcome:)